OpenStudy (anonymous):

Graph the original line and its inverse. 1. f(x)=5x+1 2. f(x)= x+8/4 3. g(x)=2x-3/4 4. h(x)=1/2(x+2)-3

OpenStudy (anonymous):

To graph them, first off, you have to find the inverse, right? I do the first one for you as sample f(x) = 5x +1 if x =0, then y = 1 . You have (0,1) if y =0 , then x = -1/5. you have (-1/5, 0) link those point to get the graph of f(x) 

OpenStudy (anonymous):

now, find inverse of f(x) by applying "switch the role" rule, you have \[x=5f^{-1}(x) +1 \\5f^{-1}(x) = x-1\\f^{-1}(x) = \dfrac{x-1}{5}\] for this function, when x =0, y = -1/5. you have (0,-1/5) when y =0, x = 1, you have (1,0) link them you get the graph of \(f^{-1}\)